# A Noise-Robust Turn-Taking System for Real-World Dialogue Robots: A Field Experiment

### v1

@@ III-A Voice Activity Projection
TABLE I: Test loss in different noise levels
@@ IV-A Experimental Overview
For this study, we deployed service robots providing route guidance in a shopping mall (LaLaport EXPOCITY1112-1 Senribanpakukouen, Suita, Osaka, Japan) and conducted field experiments. The experiment took place over two days, from February 4 to 5, 2025, with the robots operating between 11 AM and 6 PM. The baseline method and the proposed method, which will be described later, were each operated for one day. All visitors were informed about the experiment and the associated video recording through a notification board. The study was conducted on an opt-out basis, allowing visitors who did not wish to be recorded to decline participation. This study was approved by the Research Ethics Committee of Osaka University (Reference number: R1-5-18).
@@ IV-B3 Response Generation
The system includes a GPT-4o-mini-powered generative model [31]. The generative model was used to generate Sota’s responses to user’s speech input. The system incorporated a database containing comprehensive mall information for providing route guidance. When the system detect the user’s speech completion, the system combined the dialogue history with the mall information to generate an appropriate response.
@@ IV-D Result
Each condition of the proposed and baseline systems was evaluated by 15 different visitors, respectively. Table II shows the results of the users’ subjective evaluation scores. These results were tested for homogeneity of variance, and for items where homogeneity of variance was found, a Student’s $t$-test was conducted, while for items where homogeneity of variance was not found, a Welch t-test was conducted to perform statistical testing. The results showed that the proposed model had a significantly more positive impact on the three items of Smooth conversation, Ease of use, and Fulfillment of expectations.
@@
Would recommend to others † 7.13 8.13 .190

### v2

@@ Abstract
 [^0] This paper has been accepted for presentation at IEEE/RSJ International Conference on Intelligent Robots and Systems 2025 (IROS 2025) and represents the author’s version of the work.
@@ III-A Voice Activity Projection
TABLE I: Test loss for VAP task ($L_{vap}$) in different noise levels
@@ IV-A Experimental Overview
For this study, we deployed service robots providing route guidance in a shopping mall (LaLaport EXPOCITY1112-1 Senribanpakukouen, Suita, Osaka, Japan) and conducted field experiments. The experiment took place over two days, from February 4 to 5, 2025, with the robots operating between 11 AM and 6 PM. The baseline method and the proposed method, which will be described later, were each operated for one day. All visitors were informed about the experiment and the associated video recording through a notification board. The study was conducted on an opt-out basis, allowing visitors who did not wish to be recorded to decline participation. This study was approved by the Research Ethics Committee of the University of Osaka (Reference number: R1-5-18).
@@ IV-B3 Response Generation
The system includes a GPT-4o-mini-powered generative model [31]. The generative model was used to generate Sota’s responses to user’s speech input. The system incorporated a database containing comprehensive mall information for providing route guidance. When the system detected the user’s speech completion, the system combined the dialogue history with the mall information to generate an appropriate response.
@@ IV-D Result
Each condition of the proposed and baseline systems could be evaluated by 15 different visitors, respectively. Table II shows the results of the users’ subjective evaluation scores. These results were tested for homogeneity of variance, and for items where homogeneity of variance was found, a Student’s $t$-test was conducted, while for items where homogeneity of variance was not found, a Welch t-test was conducted to perform statistical testing. The results showed that the proposed model had a significantly more positive impact on the three items of Smooth conversation, Ease of use, and Fulfillment of expectations.
@@
Would recommend to others † ‡ 7.13 8.13 .190
† Homogeneity of variance, ‡ 10-point scale (1-10)